OpenStudy (anonymous):

find the coefficient of x^6 in the expansion of(5+2x^2)^7

OpenStudy (anonymous):

my memory fails me on this - ill go and check

OpenStudy (anonymous):

It uses Mimi's Crazy theorem ;)

OpenStudy (mimi_x3):

\[t_{k} = \left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k} b^{k} \]

OpenStudy (anonymous):

Mimi, that should be \( t_{k+1} \)

OpenStudy (mimi_x3):

\[t_{6}= \left(\begin{matrix}7 \\ 6\end{matrix}\right)5\times(2x^2)^6\]

OpenStudy (mimi_x3):

Nope..its only finding the coeffient

OpenStudy (mimi_x3):

Not the independent term i think..

OpenStudy (anonymous):

You think wrong then :P

OpenStudy (anonymous):

its 7C3 5^4 2^3 = 35 * 625 * 8 = 175,000

OpenStudy (anonymous):

It's crazy right Mimi? lol

OpenStudy (anonymous):

where came from 7c3 5^4 2^3

OpenStudy (mimi_x3):

I think that im not wrong, k+1 is used for something else..cant remember..

OpenStudy (mimi_x3):

& this is not crazy, the other part is crazy ::

OpenStudy (anonymous):

because you have 2x^2 in the paraenthesis it will be the fourth term in the expansion so formula is 7C3 * 5^(7-3) *2^3

OpenStudy (anonymous):

what this teory name?

OpenStudy (mimi_x3):

Binomial theorem.

OpenStudy (mimi_x3):

wait..im wrong.

OpenStudy (anonymous):

i used the formula for the (r+1) th term of the binomial expansio (a+x)^n (r+1)th term = nC r a^(n-r) x^r

OpenStudy (anonymous):

Mimi, I am sure that the general term of \( (a+b)^n \) is \( t_{k+1} = \left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k} b^{k} \)

OpenStudy (anonymous):

yup - thats it Fool

OpenStudy (mimi_x3):

Well, in my book it says.. "Specially for the expansion of (a+b)^n , the k+1th term is...that thing..cbb typing it again